# Radiative decays of $P$-wave bottom baryons from light-cone sum rules

### v1

@@ IV.4 The $[\mathbf{6}_{F},2,1,\lambda]$ doublet
Multiplets Baryon       Mass    Difference       Decay channels   Decay width   T.R.W.    T.S.W. Yang:2020zrh
($j^{P}$) (GeV) (MeV) (keV) (keV) (MeV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Omega_{b}({1\over 2}^{-})$ $6.34\pm 0.11$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $380^{+1000}_{-~{}360}$ $850^{+2200}_{-~{}830}$ $2700$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $470^{+1200}_{-~{}470}$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Omega_{b}({1\over 2}^{-})$ $6.34\pm 0.11$ $6\pm 2$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $12^{+21}_{-12}$ $19^{+33}_{-19}$ $\sim 0$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $7.1^{+12.1}_{-~{}7.0}$
$\Omega_{b}({3\over 2}^{-})$ $6.34\pm 0.09$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $6.9^{+10.0}_{-~{}6.4}$ $12^{+17}_{-11}$ $\sim 0$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $4.8^{+7.3}_{-4.5}$
$[\mathbf{6}_{F},2,1,\lambda]$ $\Omega_{b}({3\over 2}^{-})$ $6.35\pm 0.13$ $10\pm 4$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $17^{+52}_{-17}$ $21^{+69}_{-21}$ $4.6^{+3.3}_{-1.9}$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $3.7^{+16.9}_{-~{}3.7}$
$\Omega_{b}({5\over 2}^{-})$ $6.36\pm 0.12$ $\Omega_{b}^{-}({5\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.83^{+3.45}_{-0.83}$ $0.83^{+3.45}_{-0.83}$ $2.5^{+3.5}_{-1.6}$
$[\mathbf{6}_{F},1,0,\rho]$ $\Omega_{b}({1\over 2}^{-})$ $6.32\pm 0.11$ $2\pm 1$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $60^{+130}_{-~{}55}$ $70^{+150}_{-~{}65}$ $\sim 0$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $10^{+23}_{-10}$
$\Omega_{b}({3\over 2}^{-})$ $6.32\pm 0.11$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $9.1^{+20.6}_{-~{}5.0}$ $18^{+39}_{-13}$ $\sim 0$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $8.6^{+18.8}_{-~{}7.9}$
@@ B Multiplet Baryon Mass   Splitting Partial Decay Width Total width Candidate
$\Gamma(\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{{\prime}-}\gamma$) =$18^{+22}_{-18}$ keV
$\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.22\pm 0.11$ $\Gamma(\Xi_{b}({3\over 2}^{-})\to\Xi_{b}^{*}\pi$) =$1.4^{+1.0}_{-0.9}$ MeV $1.8^{+1.1}_{-1.0}$ MeV -
@@
$\Gamma(\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$) =$10.0^{+18.7}_{-~{}9.5}$ keV

### v3

@@ IV Decays of flavor $\mathbf{6}_{F}$ baryons
Multiplets Baryon       Mass    Difference     Decay channels   Coupling Constants   Decay width   T.R.W.
($j^{P}$) (GeV) (MeV) ($\rm GeV^{-1}$) (keV) (keV)
  $[\mathbf{6}_{F},0,1,\lambda]$   $\Omega_{b}({1\over 2}^{-})$ $6.34\pm 0.11$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.20^{+0.16}_{-0.10}$ $380^{+1000}_{-~{}360}$ $850^{+2200}_{-~{}830}$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.17^{+0.13}_{-0.10}$ $470^{+1200}_{-~{}470}$
$[\mathbf{6}_{F},1,1,\lambda]$ $\Omega_{b}({1\over 2}^{-})$ $6.34\pm 0.11$ $6\pm 2$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.038^{+0.016}_{-0.014}$ $12^{+21}_{-12}$ $19^{+33}_{-19}$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.022^{+0.010}_{-0.010}$ $7.1^{+12.1}_{-~{}7.0}$
$\Omega_{b}({3\over 2}^{-})$ $6.34\pm 0.09$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.027^{+0.011}_{-0.011}$ $6.9^{+10.0}_{-~{}6.4}$ $12^{+17}_{-11}$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.033^{+0.012}_{-0.013}$ $4.8^{+7.3}_{-4.5}$
$[\mathbf{6}_{F},2,1,\lambda]$ $\Omega_{b}({3\over 2}^{-})$ $6.35\pm 0.13$ $10\pm 4$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.10^{+0.15}_{-0.10}$ $17^{+52}_{-17}$ $21^{+69}_{-21}$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.055^{+0.069}_{-0.027}$ $3.7^{+16.9}_{-~{}3.7}$
$\Omega_{b}({5\over 2}^{-})$ $6.36\pm 0.12$ $\Omega_{b}^{-}({5\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.015^{+0.019}_{-0.010}$ $0.83^{+3.45}_{-0.83}$ $0.83^{+3.45}_{-0.83}$
$[\mathbf{6}_{F},1,0,\rho]$ $\Omega_{b}({1\over 2}^{-})$ $6.32\pm 0.11$ $2\pm 1$ $\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.18^{+0.10}_{-0.05}$ $60^{+130}_{-~{}55}$ $70^{+150}_{-~{}65}$
$\Omega_{b}^{-}({1\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.057^{+0.031}_{-0.017}$ $10^{+23}_{-10}$
$\Omega_{b}({3\over 2}^{-})$ $6.32\pm 0.11$ $\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{-}\gamma$ $0.070^{+0.041}_{-0.022}$ $9.1^{+20.6}_{-~{}5.0}$ $18^{+39}_{-13}$
$\Omega_{b}^{-}({3\over 2}^{-})\to\Omega_{b}^{*-}\gamma$ $0.010^{+0.05}_{-0.02}$ $8.6^{+18.8}_{-~{}7.9}$
@@ B Multiplet Baryon Mass   Splitting Partial Decay Width Total width Candidate
$\Gamma(\Xi_{b}^{\prime-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$) =$130^{+260}_{-110}$ keV
$\Xi_{b}^{\prime}({3\over 2}^{-})$ $6.22\pm 0.11$ $\Gamma(\Xi_{b}({3\over 2}^{-})\to\Xi_{b}^{*}\pi$) =$1.4^{+1.0}_{-0.9}$ MeV $1.8^{+1.4}_{-1.1}$ MeV -
@@
$\Gamma(\Xi_{b}^{\prime-}({3\over 2}^{-})\to\Xi_{b}^{-}\gamma$) =$51^{+93}_{-51}$ keV
@@ Acknowledgments
Appendix A Light-cone sum rule method and photon distribution amplitudes
The QCD sum rule method is based on the fundamental QCD Lagrangian, and it takes into account the non-perturbative nature of the QCD vacuum. In the light-cone sum rule method, a light-cone variant is further introduced to conduct the operator product expansion based on the twists of the operators, and all the non-perturbative influences are integrated into the matrix elements of non-local operators  Braun:1988qv ; Chernyak:1990ag ; Ball:1998je ; Ball:2006wn ; Ball:2004rg ; Ball:1998kk ; Ball:1998sk ; Ball:1998ff ; Ball:2007rt ; Ball:2007zt ; Wang:2007mc ; Wang:2009hra ; Aliev:2010uy ; Sun:2010nv ; Khodjamirian:2011jp ; Han:2013zg ; Offen:2013nma ; Meissner:2013hya . This method has been extensively applied in various areas of hadron physics.
The first step to calculate the coupling constants is to write the proper correlation function in terms of heavy baryon interpolating current . That is
$\Pi_{\mu\nu}(p,q)=i\int d^{4}xe^{ipx}\left<{\cal\gamma}(q)|{\cal T}\left\{J_{\mu}(x)\bar{J}_{\nu}(0)\right\}|0\right>,$ (26)
where ${\cal\gamma}(q)$ represents the photon carrying the four-momentum $q$, and $p$ represents the outgoing heavy baryon four-momentum. The above correlation function can be calculated in two ways:
  • $\bullet$
    At the hadronic level, we can calculate the correlation function by inserting the complete set of hadronic states with the same quantum numbers of the corresponding heavy baryons. It is calculated in the timelike region and contains observables like coupling constants.
  • $\bullet$
    At the quark-gluon level, we can calculate the correlation function by utilizing the method of operator product expansion in terms of light-cone distribution amplitudes of the on-shell mesons and other QCD degrees of freedom.
These two sides are matched via a dispersion integral which leads to the sum rules for the corresponding coupling constants. To suppress the contributions of the higher states and continuum, Borel transformation and continuum subtraction procedures are applied.
There are many photon distribution amplitudes used in the present study. We list them in Eqs. (27-A): $\phi_{\gamma}$ is the leading twist-2 distribution amplitude; $\psi^{(v)}$, $\psi^{(a)}$, ${\cal A}$, and ${\cal V}$ are the twist-3 ones; $h_{\gamma}$, $S$, $\widetilde{S}$, and $T_{1,2,3,4}$ are the twist-4 ones. We refer to Ref. Ball:2002ps for the detailed expressions of these photon distribution amplitudes.
[…]
$\langle 0|\bar{q}(z)g\widetilde{G}_{\mu\nu}(vz)\gamma_{\alpha}\gamma_{5}q(-z)|\gamma^{(\lambda)}(q)\rangle$ $=$ $e_{q}\,f_{3\gamma}\,q_{\alpha}[q_{\nu}e^{(\lambda)}_{\perp\mu}-q_{\mu}e^{(\lambda)}_{\perp\nu}]\int\mathcal{D}\underline{\alpha}{\cal A}(\underline{\alpha})e^{-iqz\alpha_{v}}\,,$ (30)
$\langle 0|\bar{q}(z)gG_{\mu\nu}(vz)i\gamma_{\alpha}q(-z)|\gamma^{(\lambda)}(q)\rangle$ $=$ $e_{q}\,f_{3\gamma}\,q_{\alpha}[q_{\nu}e^{(\lambda)}_{\perp\mu}-q_{\mu}e^{(\lambda)}_{\perp\nu}]\int\mathcal{D}\underline{\alpha}{\cal V}(\underline{\alpha})e^{-iqz\alpha_{v}}\,,$ (31)
$\langle 0|\bar{q}(z)g{G}_{\mu\nu}(vz)q(-z)|\gamma^{(\lambda)}(q)\rangle$ $=$ $ie_{q}\,\langle\bar{q}q\rangle[q_{\nu}e^{(\lambda)}_{\perp\mu}-q_{\mu}e^{(\lambda)}_{\perp\nu}]\int\mathcal{D}\underline{\alpha}S(\underline{\alpha})e^{-iqz\alpha_{v}}\,,$ (32)
$\langle 0|\bar{q}(z)g\tilde{G}_{\mu\nu}(vz)i\gamma_{5}q(-z)|\gamma^{(\lambda)}(q)\rangle$ $=$ $ie_{q}\,\langle\bar{q}q\rangle[q_{\nu}e^{(\lambda)}_{\perp\mu}-q_{\mu}e^{(\lambda)}_{\perp\nu}]\int\mathcal{D}\underline{\alpha}\widetilde{S}(\underline{\alpha})e^{-iqz\alpha_{v}}\,,$ (33)
$\langle 0|\bar{q}(z)\sigma_{\alpha\beta}g{G}_{\mu\nu}(vz)q(-z)|\gamma^{(\lambda)}(q)\rangle$ $=$ $e_{q}\,\langle\bar{q}q\rangle[q_{\alpha}e^{(\lambda)}_{\perp\mu}g^{\perp}_{\beta\nu}-q_{\beta}e^{(\lambda)}_{\perp\mu}g^{\perp}_{\alpha\nu}-q_{\alpha}e^{(\lambda)}_{\perp\nu}g^{\perp}_{\beta\mu}+q_{\beta}e^{(\lambda)}_{\perp\nu}g^{\perp}_{\alpha\mu}]T_{1}(v,qz)$
$+e_{q}\,\langle\bar{q}q\rangle[q_{\mu}e^{(\lambda)}_{\perp\alpha}g^{\perp}_{\beta\nu}-q_{\mu}e^{(\lambda)}_{\perp\beta}g^{\perp}_{\alpha\nu}-q_{\nu}e^{(\lambda)}_{\perp\alpha}g^{\perp}_{\beta\mu}+q_{\nu}e^{(\lambda)}_{\perp\beta}g^{\perp}_{\alpha\mu}]T_{2}(v,qz)$
$+\frac{e_{q}\,\langle\bar{q}q\rangle}{qz}[q_{\alpha}q_{\mu}e^{(\lambda)}_{\perp\beta}z_{\nu}-q_{\beta}q_{\mu}e^{(\lambda)}_{\perp\alpha}z_{\nu}-q_{\alpha}q_{\nu}e^{(\lambda)}_{\perp\beta}z_{\mu}+q_{\beta}q_{\nu}e^{(\lambda)}_{\perp\alpha}z_{\mu}]T_{3}(v,qz)$
$+\frac{e_{q}\,\langle\bar{q}q\rangle}{qz}[q_{\alpha}q_{\mu}e^{(\lambda)}_{\perp\nu}z_{\beta}-q_{\beta}q_{\mu}e^{(\lambda)}_{\perp\nu}z_{\alpha}-q_{\alpha}q_{\nu}e^{(\lambda)}_{\perp\mu}z_{\beta}+q_{\beta}q_{\nu}e^{(\lambda)}_{\perp\mu}z_{\alpha}]T_{4}(v,qz)\,.$
Appendix B Discussion on the light cone dominance
In this appendix we investigate the light cone dominance. For a detailed discussion, we refer to Refs. Mazzanti:1972zi ; Ellis:1973pf ; Frishman:1973pp ; Konetschny:1973fj ; Ellis:1971eu , and especially the doctoral dissertation of Long-Fei Gan Gan:2009zb .
To demonstrate why large momentum implies that the main contribution to the correlation function comes from the region near the light-cone, where $x^{2}$ = 0, we consider the following two-point correlation function:
$\Pi^{\mu}(\omega,\omega^{\prime})=\int d^{4}x\,e^{-ik\cdot x}\langle 0|J_{\Xi_{b}^{0}[\frac{1}{2}^{-}],1,1,\rho}(0)\bar{J}^{\mu}_{\Xi_{b}^{*}0}(x)|\gamma(q)\rangle\,.$ (35)
To derive the LCSR for this process, it is necessary to calculate the correlation function in deep Euclidean space, i.e., in the region of large $Q^{2}$ and $|(k-q)^{2}|$, and then relate the computed results to the hadronic representation through dispersion relations. Below, we explain why large momentum implies that the main contribution to the correlation function arises from the region near the light-cone, when both $Q^{2}$ and $|(k-q)^{2}|$ are large. To facilitate this, we define the variable $\nu=q\cdot k=(k^{2}-(k-q)^{2})/2$, which can also be large:
$|\nu|\sim|(k-q)^{2}|\sim Q^{2}\gg\Lambda_{QCD}^{2}\,.$ (36)
For convenience, we define the variable $\xi=2\nu/Q^{2}$, with $\xi\sim 1$ in the momentum region given by Eq. (36). Consider a reference frame in which the momentum of the photon is small compared to the momentum of the ground state baryon. i.e., $|\vec{q}|\sim\mu$, $q_{0}\sim\mu$, and $\mu^{2}<<Q^{2}$. In this reference frame, we have $k_{0}\sim Q^{2}\xi/(4\mu)+\mathcal{O}(\mu)$. Therefore, the variable $k\cdot x$ in the exponential function of the correlation function, as shown in Eq. (35), can be approximated as:
$k\cdot x$ $=k_{0}x_{0}-k_{3}x_{3}$ (37)
$\simeq\frac{Q^{2}\xi}{4\mu}x_{0}-\left(\sqrt{\frac{Q^{4}\xi^{2}}{16\mu^{2}}+Q^{2}}\right)x_{3}\simeq\frac{Q^{2}\xi}{4\mu}(x_{0}-x_{3})-\frac{2\mu}{\xi}x_{3}\,.$ (38)
To avoid drastic oscillations in the integrand of the correlation function in Eq. (35), the following conditions must be satisfied: $x_{0}-x_{3}=4\mu/(Q^{2}\xi)$ and $x_{3}=\xi/(2\mu)$. From these two conditions, we can derive the following result:
$x_{0}^{2}\simeq\left(x_{3}+\frac{4\mu}{Q^{2}\xi}\right)^{2}\simeq x_{3}^{2}+\frac{4}{Q^{2}}+O\left(\frac{\mu^{2}}{Q^{4}}\right)\,.$ (39)
This indicates that in the momentum region given by Eq. (36), $x^{2}\sim 1/Q^{2}\to 0$, which explains why the main contribution to the correlation function comes from the region near the light-cone.